OpenStudy (anonymous):

Help please will medal! Simplify (tan^2 theta + 1)/(tan^2 theta)

OpenStudy (anonymous):

explain how you solved it too!

OpenStudy (xapproachesinfinity):

hmm you have \(\frac{\tan^2\theta+1}{\tan^2x}\) remember that 1+tan^2x=sec^2x

OpenStudy (xapproachesinfinity):

so you have sec^2x/tan^2x following

OpenStudy (anonymous):

The answer is apparently csc^2 theta but I don't know how to get to that

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

asked are you following what i did! did you understand it

OpenStudy (anonymous):

not really...

OpenStudy (xapproachesinfinity):

eh sorry i used x instead of theta

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

i said that tan^2x+1=sec^2x this an identity you should know

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

so the top becomes sec^2x so we got sec^2x/tan^2x

OpenStudy (xapproachesinfinity):

good so far

OpenStudy (anonymous):

yep

OpenStudy (xapproachesinfinity):

ok, now that tanx in sin and cosine

OpenStudy (xapproachesinfinity):

what is*

OpenStudy (xapproachesinfinity):

we know that tanx=sinx/cosx correct

OpenStudy (anonymous):

correct

OpenStudy (xapproachesinfinity):

ok so it follow that tan^2x=sin^2x/cos^2x

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

now we got sec^2x/(sin^2x/cos^2x)

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

i just replaced tan^2x by what's in the bottom

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

now you flip the fraction in the bottom to get sec^2x (cos^2x/sin^2x)

OpenStudy (anonymous):

that makes sense

OpenStudy (xapproachesinfinity):

to go from division to multiplication you flip the fraction right

OpenStudy (xapproachesinfinity):

ok good

OpenStudy (anonymous):

right

OpenStudy (xapproachesinfinity):

now what is sec^2x

OpenStudy (xapproachesinfinity):

it should be 1/cos^2x since secx=1/cosx

OpenStudy (xapproachesinfinity):

right?

OpenStudy (anonymous):

right

OpenStudy (xapproachesinfinity):

okay now we got (1/cos^2x)*(cos^2x/sin^2x)=1/sin^2x after canceling cos^2x

OpenStudy (anonymous):

what!

OpenStudy (xapproachesinfinity):

you got the cancellation part?

OpenStudy (anonymous):

yeah

OpenStudy (xapproachesinfinity):

we have \(\huge \frac{1}{cos^2x}\frac{cos^2x}{sin^2x}\)

OpenStudy (anonymous):

correct

OpenStudy (xapproachesinfinity):

if we cancell we get \(\huge \frac{1}{sin^2x}\) and what it that?

OpenStudy (xapproachesinfinity):

any chance you know what that might be?

OpenStudy (anonymous):

csc^2x?

OpenStudy (xapproachesinfinity):

yeah! we have sinx=1/cscx so it follows that sin^2x=1/csc^2x

OpenStudy (anonymous):

Thank you so much for the help!!

OpenStudy (xapproachesinfinity):

that means also csc^2x=1/sin^2x it is the same thing

OpenStudy (xapproachesinfinity):

welcome!